# Notions of rank and independence in countably categorical theories

### v1

@@ 3. $n$-rank
Proof of the claim. Suppose that $a^{\prime}\in\mathrm{acl}^{n}(A)\setminus\mathrm{acl}^{n}(B)$ and $\mathrm{rk}^{n}(a^{\prime}/B)=\alpha+1$ where $\alpha\geq 1$. By the definition of $\mathrm{rk}^{n}$ there is $a\in\mathrm{acl}^{n}(a^{\prime})\setminus\mathrm{acl}^{n}(B)$ such that $\mathrm{rk}^{n}(a^{\prime}/aB)=\alpha$. If $a^{\prime}\in\mathrm{acl}^{n}(aB)$ then $\mathrm{rk}^{n}(a^{\prime}/aB)=0$, so $\alpha=0$, which contradicts that $\alpha\geq 1$. Hence $a^{\prime}\notin\mathrm{acl}^{n}(aB)$. Let $\beta=\mathrm{rk}^{n}(a/B)$, so $\beta<\omega$ and (by Lemma 3.5) there are $a_{1},\ldots,a_{\beta}\in\mathrm{acl}^{n}(a)$ such that, for all $i=1,\ldots,\beta$, $a_{i}\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{i-1}\}\cup B)$. Since $a_{1},\ldots,a_{\beta}\in\mathrm{acl}^{n}(a)\subseteq\mathrm{acl}^{n}(a^{\prime})$ and $a^{\prime}\notin\mathrm{acl}^{n}(aB)$ it follows that $a\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{\beta}\}\cup B)$. Hence the sequence $a_{1},\ldots,a_{\beta},a$ witnesses that $\mathrm{rk}^{n}(a^{\prime}/B)\geq\beta+1>\beta=\mathrm{rk}^{n}(a/B)$. As $\mathrm{rk}^{n}(a^{\prime}/B)=\alpha+1$ we must have $\beta\leq\alpha$, so $\mathrm{rk}^{n}(a/B)\leq\alpha$. As $a\notin\mathrm{acl}^{n}(B)$ we also have $\mathrm{rk}^{n}(a/B)\geq 1$. This concludes the proof of the claim.
@@
Suppose that $\bar{d}$ is a finite sequence of elements from $M_{n}$, $B,C\subseteq M^{\mathrm{eq}}$ are finite and $B\subseteq C$. Let $\alpha=\mathrm{rk}^{n}(A/B)$ (so $\alpha<\omega$). Then there is a finite sequence $\bar{d}^{\prime}$ of elements from $M_{n}$ such that $\mathrm{tp}(\bar{d}^{\prime}/B)=\mathrm{tp}(\bar{d}/B)$ and $\mathrm{rk}^{n}(\bar{d}^{\prime}/C)=\alpha$.
@@
In particular, $\mathrm{rk}^{n}(a_{1},\ldots,a_{\alpha}/B)=\alpha$ and $\mathrm{rk}^{n}(\bar{d}/\{a_{1},\ldots,a_{\alpha}\}\cup B)=0$, so (as $\mathrm{rng}(\bar{d})\subseteq M_{n}$) $\mathrm{rng}(\bar{d})\subseteq\mathrm{acl}^{n}(A)\subseteq\mathrm{acl}^{n}(\{a_{1},\ldots,a_{\alpha}\}\cup B)$ and hence $\mathrm{acl}^{n}(\bar{d}B)=\mathrm{acl}^{n}(\{a_{1},\ldots,a_{\alpha}\}\cup B)$.
@@ 4. $n$-independence
Now we are ready to define a hierarchy of independence relations, the notions of $n$-independence, for all $n<\omega$. In this section we will prove that, for all $\omega, nn-$independence has all properties of an independence relation restricted to $M_{n}$ with the possible exception of the symmetry property. Actually the restriction to $M_{n}$ is only needed for the extension property. We also show that if the theory $T$ has soft elimination of imaginaries, then the hierarchy of $n$-independence relations collapses to the bottom level of 0-independence.
@@
Let $T$ be as in Example 3.3, $\mathcal{M}\models T$ and let $a,b\in M$ be distinct and such that $a]_{E}=[b]_{E}$. By similar arguments as in that example we get $\mathrm{rk}^{0}(a/b)=\mathrm{rk}^{0}(a)$ and $\mathrm{rk}^{1}(a/b)=1<2=\mathrm{rk}^{1}(a)$. Hence $a\underset{\emptyset}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}b$ and $a\underset{\emptyset}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{1}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b$.
@@
(Transitivity) Suppose that $A,B,C,D\subseteq M^{\mathrm{eq}}$ and $B\subseteq C\subseteq D$. Then $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$ if and only if $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$ and $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$.
Proof. Suppose that $A,B,C,D\subseteq M^{\mathrm{eq}}$ and $B\subseteq C\subseteq D$. By the definition of $n$-independence it suffices to prove that for all finite $A^{\prime}\subseteq A$, $A^{\prime}\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$ if and only if $A^{\prime}\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$ and $A^{\prime}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$. Therefore we can assume that $A$ is finite. It follows from Lemma 4.4 that if $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$ then $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$ and $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$.
So suppose that $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$ and $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$. Then $\mathrm{rk}^{n}(A/B)=\mathrm{rk}^{n}(A/C)=\mathrm{rk}^{n}(A/D)$, so $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$. $\square$
@@
[…]
@@
(Extension) Let $\bar{a}$ be a finite sequence of elements from $M_{n}$, and let $B,C\subseteq M^{\mathrm{eq}}$ where $B\subseteq C$. Then there is $\bar{d}^{\prime}\subseteq M^{\mathrm{eq}}$ such that $\mathrm{tp}(\bar{d}^{\prime}/B)=\mathrm{tp}(\bar{d}/B)$ and $\bar{d}^{\prime}\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$.
@@
Suppose that $T$ has soft elimination of imaginaries. For all $n<\omega$ and all $A,B,C\subseteq M^{\mathrm{eq}}$ we have $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}B$.
Proof. By definition of $\smile$$|^{n}$, $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $\mathrm{rk}^{n}(A^{\prime}/BC)=\mathrm{rk}^{n}(A^{\prime}/C)$ for all finite $A^{\prime}\subseteq A$. Under the assumption that $T$ has soft elimination of imaginaries, Lemma 3.8 implies that $\mathrm{rk}^{n}(A^{\prime}/BC)=\mathrm{rk}^{0}(A^{\prime}/BC)$ and $\mathrm{rk}^{n}(A^{\prime}/C)=\mathrm{rk}^{0}(A^{\prime}/C)$ for all finite $A^{\prime}\subseteq A$. Hence $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}B$. $\square$
@@ 5. The exchange property and symmetry
Now suppose that, for some $m$, $k_{1}<\ldots<k_{m}\leq\alpha$ and $a_{1},\ldots,a_{k_{m}}\in\mathrm{acl}^{n}(A)$ have been defined in such a way that, for all $j=0,\ldots,m-1$, $\{a_{k_{j}+1},\ldots,a_{k_{j+1}}\}$ is a basis of
@@
Suppose that $A\not\subseteq\mathrm{acl}^{n}(\{a_{1},\ldots,a_{k_{m}}\}\cup B)$. By Lemma 3.10 the set
@@
be a basis of $X_{m+1}$. Note that $k_{m+1}$ depends only on $k_{m}$ and the dimension of $X_{m+1}$ where the later depends only on $A$, $B$, $a_{1},\ldots,a_{k_{m}}$ and $n$. The inductive assumption that for any choice of $a^{\prime}_{1},\ldots,a^{\prime}_{k_{m}}$ with the same properties as $a_{1},\ldots,a_{k_{m}}$ we have $\mathrm{acl}^{\mathrm{eq}}(\{a^{\prime}_{1},\ldots,a^{\prime}_{k_{m}}\}\cup B)=\mathrm{acl}^{\mathrm{eq}}(\{a_{1},\ldots,a_{k_{m}}\}\cup B)$ implies that $\mathrm{acl}^{\mathrm{eq}}(\{a^{\prime}_{1},\ldots,a^{\prime}_{k_{m+1}}\}\cup B)=\mathrm{acl}^{\mathrm{eq}}(\{a_{1},\ldots,a_{k_{m+1}}\}\cup B)$ for any basis $\{a^{\prime}_{k_{m}+1},\ldots,a^{\prime}_{k_{m+1}}\}\subseteq X_{m+1}$. It also implies that, ultimately, $k_{m+1}$ depends only on $A$, $B$ and $n$.
@@
Proof. Suppose that $\mathrm{rk}n(A/B)=\alpha$. By Lemma 5.4, it suffices to prove that there is at least one $n$-$c$s for $A/B$ that has length $\alpha$. By Lemma 3.5 and Definition 3.6 of $n$-cs, there is an $n$-cs $a_{1},\ldots,a_{\alpha}$ for $A/B$. By the same definition and lemma we have
(5.1) $a_{i}\notin\mathrm{acl}^{n}(a_{i}/\{a_{1},\ldots,a_{i-1}\}\cup B)\ \text{ and }$
@@
Suppose that, for some $0\leq k\leq\alpha$ and $1\leq s\leq\alpha-k$, we have
@@
By Lemmas 5.4 and 5.6, there is an nnccs $b^{\prime}_{1},\ldots,b^{\prime}_{\beta}$ for $B/AC$ with core sequence of indices $0=l_{0}<l_{1}<\ldots<l_{\gamma}=\beta$. By part (i), $b^{\prime}_{1},\ldots,b^{\prime}_{\beta}$ is also an $n$-$c$s for $B/C$ with the same core sequence of indices $0=l_{0}<l_{1}<\ldots<l_{\gamma}=\beta$. By Lemma 5.4, we then have $\gamma=\delta$ and $l_{m}=k_{m}$ for all $m=0,\ldots,\delta$. The same lemma also tells that, for all $m=1,\ldots,\delta$,
(5.6) $\mathrm{acl}^{\mathrm{eq}}(\{b_{1},\ldots,b_{k_{m}}\}\cup C)=\mathrm{acl}^{\mathrm{eq}}(\{b^{\prime}_{1},\ldots,b^{\prime}_{k_{m}}\}\cup C)$
and hence

### v2

@@ 3. $n$-rank
Proof of the claim. Suppose that $a^{\prime}\in\mathrm{acl}^{n}(A)\setminus\mathrm{acl}^{n}(B)$ and $\mathrm{rk}^{n}(a^{\prime}/B)=\alpha+1$ where $\alpha\geq 1$. Suppose that $a^{\prime}\in\mathrm{acl}^{n}(A)\setminus\mathrm{acl}^{n}(B)$ and $\mathrm{rk}^{n}(a^{\prime}/B)=\alpha+1$ where $\alpha\geq 1$. By the definition of $\mathrm{rk}^{n}$ there is $a\in\mathrm{acl}^{n}(a^{\prime})\setminus\mathrm{acl}^{n}(B)$ such that $\mathrm{rk}^{n}(a^{\prime}/aB)=\alpha$. If $a^{\prime}\in\mathrm{acl}^{n}(aB)$ then $\mathrm{rk}^{n}(a^{\prime}/aB)=0$, so $\alpha=0$, which contradicts that $\alpha\geq 1$. Hence $a^{\prime}\notin\mathrm{acl}^{n}(aB)$. As $a\notin\mathrm{acl}^{n}(B)$ we have $\mathrm{rk}^{n}(a/B)\geq 1$.
We now show that $\mathrm{rk}^{n}(a/B)\leq\alpha$. For a contradiction, suppose that $\mathrm{rk}^{n}(a/B)\geq\alpha+1$. By Lemma 3.5 there are $a_{1},\ldots,a_{\alpha+1}\in\mathrm{acl}^{n}(a)$ such that for all $i=1,\ldots,\alpha+1$, $a_{i}\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{i-1}\}\cup B)$. If $a^{\prime}\in\mathrm{acl}^{n}(\{a_{1},\ldots,a_{\alpha+1}\}\cup B)$, then $a^{\prime}\in\mathrm{acl}^{n}(aB)$ contradicting what we concluded above. Hence $a^{\prime}\notin\mathrm{acl}^{n}(aB)$. Therefore the sequence $a_{1},\ldots,a_{\alpha+1},a^{\prime}\in\mathrm{acl}^{n}(a^{\prime})$ and Lemma 3.5 witness that $\mathrm{rk}^{n}(a^{\prime}/B)\geq\alpha+2$ which contradicts the assumption about $a^{\prime}$. This concludes the proof of the claim.
@@
Suppose that $\bar{d}$ is a finite sequence of elements from $M_{n}$, $B,C\subseteq M^{\mathrm{eq}}$ are finite and $B\subseteq C$. Let $\alpha=\mathrm{rk}^{n}(\bar{d}/B)$ (so $\alpha<\omega$). Then there is a finite sequence $\bar{d}^{\prime}$ of elements from $M_{n}$ such that $\mathrm{tp}(\bar{d}^{\prime}/B)=\mathrm{tp}(\bar{d}/B)$ and $\mathrm{rk}^{n}(\bar{d}^{\prime}/C)=\alpha$.
@@
In particular, $\mathrm{rk}^{n}(a_{1},\ldots,a_{\alpha}/B)=\alpha$ and $\mathrm{rk}^{n}(\bar{d}/\{a_{1},\ldots,a_{\alpha}\}\cup B)=0$, so (as $\mathrm{rng}(\bar{d})\subseteq M_{n}$) $\mathrm{rng}(\bar{d})\subseteq\mathrm{acl}^{n}(\bar{d})\subseteq\mathrm{acl}^{n}(\{a_{1},\ldots,a_{\alpha}\}\cup B)$ and hence $\mathrm{acl}^{n}(\bar{d}B)=\mathrm{acl}^{n}(\{a_{1},\ldots,a_{\alpha}\}\cup B)$.
@@ 4. $n$-independence
Now we are ready to define a hierarchy of independence relations, the notions of $n$-independence, for all $n<\omega$. In this section we will prove that, for all $\omega, nn-$independence has all properties of an independence relation restricted to $M_{n}$ with the possible exception of the symmetry property. Actually the restriction to $M_{n}$ is only needed for the extension property. We also show that if the theory $T$ has geometric elimination of imaginaries, then the hierarchy of $n$-independence relations collapses to the bottom level of 0-independence.
@@
Let $T$ be as in Example 3.3, $\mathcal{M}\models T$ and let $a,b\in M$ be distinct and such that $a]_{E}=[b]_{E}$. By the conclusions in that example we get $\mathrm{rk}^{0}(a/b)=\mathrm{rk}^{0}(a)$ and $\mathrm{rk}^{1}(a/b)=1<2=\mathrm{rk}^{1}(a)$. Hence $a\underset{\emptyset}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}b$ and $a\underset{\emptyset}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{1}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}b$.
@@
(Transitivity) Suppose that $A,B,C,D\subseteq M^{\mathrm{eq}}$ and $B\subseteq C\subseteq D$. If $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$ and $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$ then $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$.
Proof. Suppose that $A,B,C,D\subseteq M^{\mathrm{eq}}$ and $B\subseteq C\subseteq D$. By the definition of $n$-independence it suffices to prove that for all finite $A^{\prime}\subseteq A$, if $A^{\prime}\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$ and $A^{\prime}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$ then $A^{\prime}\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$. Therefore we can assume that $A$ is finite. So suppose that $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$ and $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$. Then $\mathrm{rk}^{n}(A/B)=\mathrm{rk}^{n}(A/C)=\mathrm{rk}^{n}(A/D)$, so $A\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}D$. $\square$
@@
Proof. If $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}B$ then (by definition of $\smile$$|^{n}\diagup$) $A^{\prime}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}B$ for some finite $A^{\prime}\subseteq A$. Then $\mathrm{rk}^{n}(A^{\prime}/BC)<\mathrm{rk}^{n}(A^{\prime}/C)$. By Lemma 3.11 (i) there is finite $B^{\prime}\subseteq B$ such that $\mathrm{rk}^{n}(A^{\prime}/B^{\prime}C)$ << $\mathrm{rk}^{n}(A^{\prime}/C)$. Hence $A^{\prime}\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}\hskip-12.0pt\raisebox{3.0pt}[5.0pt][0.0pt]{$\diagup$}}B^{\prime}$. $\square$
@@
(Extension) Let $\bar{a}$ be a finite sequence of elements from $M_{n}$, and let $B,C\subseteq M^{\mathrm{eq}}$ where $B\subseteq C$. Then there is $\bar{a}^{\prime}\subseteq M^{\mathrm{eq}}$ such that $\mathrm{tp}(\bar{a}^{\prime}/B)=\mathrm{tp}(\bar{a}/B)$ and $\bar{a}^{\prime}\underset{B}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}C$.
@@
Suppose that $T$ has geometric elimination of imaginaries. For all $n<\omega$ and all $A,B,C\subseteq M^{\mathrm{eq}}$ we have $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}B$.
Proof. By definition of $\smile$$|^{n}$, $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $\mathrm{rk}^{n}(A^{\prime}/BC)=\mathrm{rk}^{n}(A^{\prime}/C)$ for all finite $A^{\prime}\subseteq A$. Under the assumption that $T$ has geometric elimination of imaginaries, Lemma 3.8 implies that $\mathrm{rk}^{n}(A^{\prime}/BC)=\mathrm{rk}^{0}(A^{\prime}/BC)$ and $\mathrm{rk}^{n}(A^{\prime}/C)=\mathrm{rk}^{0}(A^{\prime}/C)$ for all finite $A^{\prime}\subseteq A$. Hence $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}B$. $\square$
@@ 5. The exchange property and symmetry
Now suppose that, for some $m$, $k_{1}<\ldots<k_{m}$ and $a_{1},\ldots,a_{k_{m}}\in\mathrm{acl}^{n}(A)$ have been defined in such a way that, for all $j=0,\ldots,m-1$, $\{a_{k_{j}+1},\ldots,a_{k_{j+1}}\}$ is a basis of
@@
Suppose that $\mathrm{acl}^{n}(A)\not\subseteq\mathrm{acl}^{n}(\{a_{1},\ldots,a_{k_{m}}\}\cup B)$. By Lemma 3.10 the set
@@
be a basis of $X_{m+1}$. Note that $k_{m+1}$ depends only on $k_{m}$ and the dimension of $X_{m+1}$ where the latter depends only on $A$, $B$, $a_{1},\ldots,a_{k_{m}}$ and $n$. The inductive assumption that for any choice of $a^{\prime}_{1},\ldots,a^{\prime}_{k_{m}}$ with the same properties as $a_{1},\ldots,a_{k_{m}}$ we have $\mathrm{acl}^{\mathrm{eq}}(\{a^{\prime}_{1},\ldots,a^{\prime}_{k_{m}}\}\cup B)=\mathrm{acl}^{\mathrm{eq}}(\{a_{1},\ldots,a_{k_{m}}\}\cup B)$ implies that $\mathrm{acl}^{\mathrm{eq}}(\{a^{\prime}_{1},\ldots,a^{\prime}_{k_{m+1}}\}\cup B)=\mathrm{acl}^{\mathrm{eq}}(\{a_{1},\ldots,a_{k_{m+1}}\}\cup B)$ for any basis $\{a^{\prime}_{k_{m}+1},\ldots,a^{\prime}_{k_{m+1}}\}\subseteq X_{m+1}$. It also implies that, ultimately, $k_{m+1}$ depends only on $A$, $B$ and $n$.
@@
Proof. Suppose that $\mathrm{rk}^{n}(A/B)=\alpha$. By Lemma 5.4, it suffices to prove that there is at least one $n$-$c$s for $A/B$ that has length $\alpha$. By Lemma 3.5 and Definition 3.6 of $n$-cs, there is an $n$-cs $a_{1},\ldots,a_{\alpha}$ for $A/B$. By the same definition and lemma we have
(5.1) $a_{i}\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{i-1}\}\cup B)\ \text{ and }$
@@
Suppose that, for some $0\leq k<\alpha$ and $1\leq s\leq\alpha-k$, we have
@@
By Lemmas 5.4 and 5.6, there is an $n$-$c$s $b^{\prime}_{1},\ldots,b^{\prime}_{\beta}$ for $B/AC$ with core sequence of indices $0=l_{0}<l_{1}<\ldots<l_{\gamma}=\beta$. By part (i), $b^{\prime}_{1},\ldots,b^{\prime}_{\beta}$ is also an $n$-$c$s for $B/C$ with the same core sequence of indices $0=l_{0}<l_{1}<\ldots<l_{\gamma}=\beta$. By Lemma 5.4, we then have $\gamma=\delta$ and $l_{m}=k_{m}$ for all $m=0,\ldots,\delta$. The same lemma also tells that, for all $m=1,\ldots,\delta$,
(5.6) $\mathrm{acl}^{\mathrm{eq}}(\{b_{1},\ldots,b_{k_{m}}\}\cup C)=\mathrm{acl}^{\mathrm{eq}}(\{b^{\prime}_{1},\ldots,b^{\prime}_{k_{m}}\}\cup C).$
Now we prove that
@@
Suppose that $d\in\mathrm{acl}^{\mathrm{eq}}(\{b_{1},\ldots,b_{k_{m}}\}\cup AC)$. Then there are a formula $\varphi(u,\bar{x},\bar{y},\bar{z})$, a tuple $\bar{b}$ of elements from $\{b_{1},\ldots,b_{k_{m}}\}$, a tuple $\bar{a}$ of elements from $A$, and a tuple $\bar{c}$ of elements from $C$, such that $\mathcal{M}^{\mathrm{eq}}\models\varphi(d,\bar{b},\bar{a},\bar{c})$ and only finitely many elements satisfy $\varphi(u,\bar{b},\bar{a},\bar{c})$. Let $m$ be such that $d$ and all elements that belong to any of $\bar{b}$, $\bar{a}$, or $\bar{c}$, are members of $M_{m}$. As $\mathcal{M}_{m}$ is $\omega$-categorical there is a formula $\theta(u,\bar{x},\bar{y},\bar{z})$ which isolates $\mathrm{tp}_{\mathcal{M}_{m}}(d,\bar{b},\bar{a},\bar{c})$, where $\mathrm{tp}_{\mathcal{M}_{m}}$’ denotes the type computed in $\mathcal{M}_{m}$. Then only finitely many elements satisfy $\theta(u,\bar{b},\bar{a},\bar{c})$, all of them belong to $M_{m}$, and $d$ is one among them. From (5.6) we have $\mathrm{rng}(\bar{b})\subseteq\mathrm{acl}^{m}(\{b^{\prime}_{1},\ldots,b^{\prime}_{k_{m}}\}\cup C)$, so there are a formula $\psi(\bar{x},\bar{y},\bar{z})$, a tuple $\bar{b}^{\prime}$ of elements from $\{b^{\prime}_{1},\ldots,b^{\prime}_{k_{m}}\}$, and a tuple $\bar{c}^{\prime}$ of elements from $C$ such that $\mathcal{M}^{\mathrm{eq}}\models\psi(\bar{b},\bar{b}^{\prime},\bar{c}^{\prime})$ and $\psi(\bar{x},\bar{b}^{\prime},\bar{c}^{\prime})$ is satisfied by only finitely many tuples. It is now clear that $d$ satisfies the formula
$\chi(u,\bar{a},\bar{b}^{\prime},\bar{c},\bar{c}^{\prime}):=\exists\bar{x}\big[\theta(u,\bar{x},\bar{a},\bar{c})\wedge\psi(\bar{x},\bar{b}^{\prime},\bar{c}^{\prime})\big].$
The formula $\psi(\bar{x},\bar{b}^{\prime},\bar{c}^{\prime})$ is satisfied by only finitely many tuples, and if $\bar{b}^{*}$ is one of them and $\mathcal{M}^{\mathrm{eq}}\models\theta(d^{\prime},\bar{b}^{*},\bar{a},\bar{c})$ for some $d^{\prime}$, then $d^{\prime}$ and all members of $\bar{b}^{*}$ belong to $M_{m}$, and $\mathrm{tp}_{\mathcal{M}_{m}}(d^{\prime},\bar{b}^{*},\bar{a},\bar{c})=\mathrm{tp}_{\mathcal{M}_{m}}(d,\bar{b},\bar{a},\bar{c})$, so only finitely many elements satisfy $\theta(u,\bar{b}^{*},\bar{a},\bar{c})$. It now follows that $\chi(u,\bar{a},\bar{b}^{\prime},\bar{c},\bar{c}^{\prime})$ is satisfied by only finitely many elements. Hence
$d\in\mathrm{acl}^{\mathrm{eq}}(\bar{a},\bar{b}^{\prime},\bar{c},\bar{c}^{\prime})\subseteq\mathrm{acl}^{\mathrm{eq}}(\{b^{\prime}_{1},\ldots,b^{\prime}_{k_{m}}\}\cup AC).$
This shows the inclusion from left to right in (5.7) and by a similar argument (letting the $b_{i}$ and $b^{\prime}_{i}$ switch roles) the converse inclusion follows. This proves (5.7).